Define and prove the Bayes’ theorem.


1
Expert's answer
2020-12-08T07:24:34-0500

Bayes’ theorem states that, if "A_1, A_2, \u2026, A_n" are mutually exclusive and exhaustive events in a sample space s, such that "P(A_i) > 0" for "i = 1, 2, \u2026, n" and E is any event which is subset of "\bigcup_{i=1}^{n}A_i \; with\; P(E) > 0" , then

"P (A_k|E) = \frac{P(A_k) \times P(E|A_k)}{\sum_{i=1}^{n}P(A_i) \times P(E|A_i)}" for "k = 1, 2, 3 \u2026 n" .

Proof:

Since "A_1, A_2,...A_n" are mutually exclusive and exhaustive events in a sample space s, we have "\bigcup_{i=1}^{n}A_i….s"

Since "A_1, A_2, A_3...A_n" are mutually disjointed, "E \cap A_1, E \cap A_2, E \cap A_3, … E \cap A_n" are mutually disjointed.

"P(E) = P(E \cap S) \\

= P[E \cap (\bigcup_{i=1}^{n}A_i)] \\

= P(\bigcup_{i=1}^{n}E \cap A_i) \\

= \sum_{i=1}^{n}P(E \cap A_i) \\

= \sum_{i=1}^{n}P(A_i) \times P(E|A_i)"

by multiplication theorem of probability.

Also we have, for k = 1, 2,….,n

"P(E \cap A_k) = P(E) \times P(A_k|E) \\

P(A_k|E) = \frac{P(E \cap A_k)}{P(E)} \\

= \frac{P(A_k) \times P(E|A_k)}{\sum_{i=1}^{n}P(A_i) \times P(E|A_i)}"